Anthracite is a type of coal that is composed of carbon with small amounts of hydrogen, oxygen, nitrogen, and sulfur. The formula for a sample of anthracite is C240H90O4NS. Calculate the amount of carbon in 55.0 kg of the sample.

mass:_______kg C
Answer #1


Molar mass of C240H90O4NS,
MM = 240*MM(C) + 90*MM(H) + 4*MM(O) + 1*MM(N) + 1*MM(S)
= 240*12.01 + 90*1.008 + 4*16.0 + 1*14.01 + 1*32.07
= 3083.2 g/mol


mass(C240H90O4NS)= 55.0 Kg
= 55000.0 g

use:
number of mol of C240H90O4NS,
n = mass of C240H90O4NS/molar mass of C240H90O4NS
=(5.5*10^4 g)/(3.083*10^3 g/mol)
= 17.84 mol
This is number of moles of C240H90O4NS

one mole of C240H90O4NS contains 240 mole of C
use:
number of moles of C = 240 * number of moles of C240H90O4NS
= 240 * 17.84
= 4.281*10^3

Molar mass of C = 12.01 g/mol

use:
mass of C,
m = number of mol * molar mass
= 4.281*10^3 mol * 12.01 g/mol
= 5.142*10^4 g
= 51.4 Kg
Answer: 51.4 Kg